OpenStudy (anonymous):

Please helpppp!!!!!!! At what interest rate the must $4500 be compounded annually to equal $10,174.07 after 14 years? Thanks! ;)

OpenStudy (anonymous):

interest rate=180

OpenStudy (anonymous):

Noo, the answer is suppose to be 6%.

OpenStudy (anonymous):

ok lol hang on.

OpenStudy (anonymous):

Can you give me the step if you don't might? Thanks!

OpenStudy (anonymous):

I did it, but I want to make sure it is the right formula..

OpenStudy (anonymous):

the formula is A=P(1+r/n)^nt a=compounded interest r=6% P=amount ($4500) n=how many times a year t=years(14)

OpenStudy (anonymous):

Nop, I still don't get it...

OpenStudy (anonymous):

I have from a similar question, this.. ln(10174.07/4500)/0.14. And i got 5.8. so it says (round it to the nearest percent.)

OpenStudy (anonymous):

@jim_thompson5910, hi life saver.. I would like to know if this is the right formula for this problem.?? ln(10174.07/4500)/0.14 Thanks! ;)

jimthompson5910 (jim_thompson5910):

A=P(1+r/n)^(nt) 10174.07=4500(1+x/1)^(1*14) 10174.07=4500(1+x)^(14) 10174.07/4500 = (1+x)^(14) 2.260904 = (1+x)^(14) (2.260904)^(1/14) = 1+x 1.06000000148 = 1+x 1.06000000148 - 1 = x 0.06000000148 = x x = 0.06000000148 x = 0.06 So the interest rate is roughly 6%

OpenStudy (anonymous):

Waooo, I am in shock.... Loved it... Thanks so much!!